# On The Implicit Large Eddy Simulation of Turbomachinery Flows Using The Flux Reconstruction Method

### v1

@@ Abstract
A high order flux reconstruction solver has been developed and validated to perform implicit large eddy simulations of industrially representative turbomachinery flows. The T106c low-pressure turbine and VKI LS89 high-pressure turbine cases are studied. The solver uses the Rusanov Riemann solver to compute the inviscid fluxes on the wall boundaries, and HLLC or Roe to evaluate inviscid fluxes for internal faces. The impact of Riemann solvers is demonstrated in terms of accuracy and non-linear stability for turbomachinery flows. It is found that HLLC is more robust than Roe, but both Riemann solvers produce very similar results if stable solutions can be obtained. For non-linear stabilization, a local modal filter, which combines a smooth indicator and a modal filter, is used to stabilize the solution. This approach requires a tuning parameter for the smoothness criterion. Detailed analysis has been provided to guide the selection of a suitable value for different spatial orders of accuracy. This local-modal filter is also compared with the recent positivity-preserving entropy filter in terms of accuracy and stability for the LS89 turbine case. The entropy filter could stabilize the computation but is more dissipative than the local modal filter. Regarding the spanwise spacing of the grid, the case of the LS89 turbine shows that a $z^{+}$ of approximately $45-60$ is suitable for obtaining a satisfactory prediction of the heat transfer coefficient of the mean flow. This would allow for a coarse grid spacing in the spanwise direction and a cost-effective ILES aerothermal simulation for turbomachinery flows.
@@
discontinuous spectral element method , high order flux reconstruction , large eddy simulation , turbomachinery
@@
[inst1]organization=Oxford Thermofluids Institute, Department of Engineering Science,addressline=University of Oxford, country=United Kingdom
@@
Effect of Riemann solvers on accuracy and numerical stability is investigated for turbomachinery flows
@@
The positivity preserving entropy filter is compared with the local modal filter and assessed in turbomachinery flows
@@ 1 Introduction
Computational Fluid Dynamics (CFD) has progressed significantly over the past few decades and transformed the design process in the aerospace industry [1]. With the advancement of computing power and numerical algorithms, scale-resolving simulations, such as Large Eddy Simulation (LES), start complementing not only the experimental campaigns but also low-order numerical tools at the design stage. Reynolds-averaged Navier-Stokes (RANS) simulations and unsteady RANS (URANS) are the current workhorse in the industry for aerothermal design. However, it is well known that the RANS/URANS approach can face difficulties in reliably predicting vortex-dominated and separated flows, and its associated heat transfer phenomenon.
The use of discontinuous spectral element methods (DSEM) has grown in prevalence in the past few years for LES due to their ability to achieve high-order spatial accuracy in unstructured grids and a compact stencil that is suited for massively parallel computations. Popular DSEM methods are the nodal Discontinuous Galerkin (DG) approach [2, 3] and, more recently, the Flux Reconstruction (FR) [4, 5] or the Correction Procedure via Reconstruction (CPR) [6] method. Despite the differences in these two methods, the numerical procedures developed in the collocation-based nodal DG could be applied to FR/CPR (such as filtering and boundary treatment) with little modification, and vice versa. Implicit LES (ILES) is frequently used in DSEM, as the built-in dissipation of DSEM is found to be adequate as a subgrid model. The superior performance of DSEM over traditional second-order finite-volume solvers on ILES has been demonstrated by Vermeire et al. [7] and Jia and Wang [8].
@@
Filtering is an attractive stabilization technique due to its ease of implementation and computational efficiency. However, care must be taken when filtering is applied to DSEM, as it can significantly undermine the accuracy of DSEM [30]. The amount of filtering should be applied as much as needed to stabilize the solution, but also as little as possible to avoid excessive dissipation. A modal filter could be localized using a smooth indicator so that the smooth-flow region will not be subject to filtering. However, this strategy requires a smoothness criterion to determine whether the element should be filtered or not. This method has not been systematically studied in the literature for turbomachinery flows, and the choice of the smooth criterion is also not clear. The first contribution of this paper is: we will demonstrate the performance of this local modal filtering approach on turbomachinery flows and perform a detailed analysis of the smoothness criterion to provide a guide to select a suitable value. The recent entropy filter of Dzanic and Witherden [31] can also be considered as a localized modal filtering approach. Physical constraints are used as a “smoothness” criterion to determine which element should be filtered. This entropy filter is implemented and its performance is assessed in turbomachinery flows and compared with the local modal filter.
Riemann solvers have been found to have an impact on accuracy and numerical stability on simple turbulent flows [13] using DSEM. These effects have not been systematically studied for turbomachinery flows. In this work, such a study will be performed on industrially representative turbomachinery cases. In addition, Alhawwary and Wang [32] performed a study of the spacing of the span of the grid on the prediction of mean flow variables (i.e. isentropic Mach number) in the T106c turbine using CPR. However, there have been no similar studies on the prediction of heat transfer coefficients in turbine blades, which is more challenging to predict than the pressure distribution on the blade surface. In this paper, we demonstrate this effect using the VKI LS89 turbine case. These are the second contribution of this work.
The computations in this paper are performed with a new in-house FR solver, AeroThermal High-Order Simulation (ATHOS). This paper is organized as follows: The governing equations of compressible flow and the FR method are briefly described. Then the local modal filter is introduced. The performance of the local modal filter on Taylor Green vortex, T106c, and LS89 turbine cases is demonstrated, the effect of Riemann solvers on accuracy and numerical stability is analyzed and a detailed analysis of the smoothness criterion of the modal filter is presented. This is then followed by the conclusions and future work.
@@ 2.2 High-Order Flux Reconstruction
The FR method is an efficient discontinuous spectral element method to solve the partial differential equation (PDE) in differential form. It is briefly described here, and for a detailed description one can refer to Huynh [4], Vicent et al. [33] and Castonguay [34]. The NS equation of the compressible flow (Equation 1) is a second order PDE. Before applying FR, it is first cast into a set of first order PDEs and this is shown in Equation 2.2 [29]:
@@ 2.3 Stabilization with filtering
DSEM methods can face robustness issues, and stabilization techniques are usually required to enhance robustness to handle underresolved flows or strong discontinuity (e.g. shocks). In this work, a local modal filtering (LMF) approach is developed. The general concept is to apply the modal filter in regions where stabilization is required while leaving the smooth-flow region untouched. Modal filtering is a simple and classical approach in DSEM and it can be represented as a matrix operation for the nodal solutions as [29]:
@@
where $\eta_{i}=\frac{i}{N_{p}-1}$, $i$ is the mode index and $N_{p}$ is the maximum order of the polynomial. $\eta_{c}$ is the cutoff mode threshold, below which the modes will not be filtered. $\alpha$ and $s$ determine the strength of the filter. $s$ is an even number, and its values are usually from 2 to 8. $s$ is related to the scale selectivity of the filter and in practice a lower value of $s$ normally leads to a stronger filter. The value of $\alpha$ is $-\log_{10}(\varepsilon_{m})$ and $\varepsilon_{m}$ is the machine zero [29]. For simulations with double precision $\alpha\approx 36$. It is obvious that the power of the exponent in Equation 18 is a negative number; therefore, the modal filter is always dissipative.
@@ 4.1 Taylor Green Vortex at Re=1600 and Re=5000
Figure 2 shows the predicted time evolution of $\varepsilon_{K}$ using LMF at Reynolds numbers of 1600 and 5000, respectively. The results are compared with DNS data. Time $t$ is the normalized time, which is computed as $t=t_{\text{physical}}/t_{c}$. $t_{c}$ is the characteristic time and it is defined as $L/V_{0}$. At Re=1600, the DNS data is from Gassner and Beck [17]. For Re=5000 it is computed by Dairay et al. [45] using the Incompact3d code. From the comparison, it can be seen that there is no noticeable difference between the results of LMF and the one without filtering. Both results yield excellent agreement with the reference DNS data. This confirms that LMF does not introduce noticeable dissipation to the simulation in the absence of strong discontinuity in the flow.
@@
The energy spectra is also examined at both Reynolds numbers and it is computed at $t=9$ and $t=14$, respectively. This is shown in Fig. 3. The spectra of the DNS data at Re=1600 is computed using a pseudo spectral method on a $512^{3}$ mesh by Carton de Wiart et al. [46] and the one for Re=5000 is again computed by Dairay et al. [45]. The maximum resolvable wavenumber is marked by a vertical dash line. It can be seen that there is excellent agreement with the DNS data and there is no noticeable difference when LMF is used or not. It is noted that the purpose of this study is to demonstrate that when the flow is smooth, LMF essentially has a negligible effect on the flow field. But for more complicated flows, LMF will be switched on and this will be demonstrated in the following two turbine cases.
@@ 4.2 T106c Low Pressure Turbine
This test case performs the spanwise periodic ILES of the transitional and separated flow on the T106c turbine cascade. This test case has been well studied in the literature and is a standard test case in the International Workshop on High-Order CFD Methods. The flow in the T106c turbine is subsonic and is calculated at Re = 80,000. The Reynolds number is based on the blade chord and the flow conditions at the exit boundary. The inlet turbulence is very low, and the flow features laminar separation and transition in the reattachment zone. The details of the geometry and boundary conditions are summarized in Table 3.
@@
The grid is generated by Gmsh [47] and the script is provided by the 4th International Workshop on High-Order CFD methods  [^1] https://how4.cenaero.be/content/as2-spanwise-periodic-dnsles-transitional-turbine-cascades. A quadrilateral mesh is generated in the blade-to-blade section and a hexahedral mesh is generated by extruding the 2D quadrilateral mesh by $10\%$ of the chord for 12 layers. The script for the “baseline” mesh from the workshop is used and the only modification to the original script is that the expansion ratio of the boundary layer mesh is reduced from 1.3 to 1.2. The total number of hexahedral cells is 122904. For a $p3$ simulation, the total number of DOF per equation is 7865856. The axial distance between the leading edge and the inlet boundary is approximately $1.6c$, where $c$ is the blade chord. The axial distance between the trailing edge and the exit boundary is approximately $2.4c$. Figure 4 shows the quadrilateral mesh in the blade-to-blade section and a close-up view of the boundary layer mesh around the trailing edge.
@@
Table 4 shows the numerical setup for the T106c turbine. LMF is used to stabilize the computation, HLLC and Roe Riemann solvers are used and compared. The values of $\alpha$ and $s$ are set to 36 and 6, respectively, and their values are fixed throughout this paper. According to Hesthaven and Warburton [29] both values provide a reasonable balance of numerical accuracy and stability. $s_{0}$ is the flow smoothness criterion and determines how ”local” the filter will be. If a large negative number is used, say $-16$, the filter will be applied to all elements. A value of -4 is used for both HLLC and Roe and successfully stabilizes the computation. More analysis of the smooth criterion can be found in Section 4.4.1.
The computation starts from a $p_{0}$ run to efficiently wash away the initial transient flow field and creates a suitable initial flow field. Then the computation is restarted and the polynomial order is increased to $p1$. The same procedures are then used to increase the order of accuracy from $p1$ to $p2$, and eventually from $p2$ to $p3$. A fixed CFL number of 0.85 is used in the calculation. For different spatial orders of accuracy and this value is scaled by $\frac{1}{2p+1}$, where $p$ is the order of polynomial. For a $p3$ computation, the average time step is approximately $0.47\times 10^{-4}t_{c}$ for HLLC-LMF, where $t_{c}$ is the characteristic time and is defined as:
@@
Figure 5 shows the isosurfaces of the Q criterion for the instantaneous flow calculated by HLLC-LMF. Transitional flows on the suction side and vortex shedding can be clearly seen. Figure 6 shows the power spectral density (PSD) of the pressure at a point downstream of the trailing edge. The point is on the middle span plane and its coordinate on the blade-to-blade section is $(0.87619-0.51794)$. The frequency is represented in terms of the Strouhal number as:
@@
in which $f$ is the frequency, $L_{char}$ is the characteristic length and $u_{char}$. In this T106c case, $L_{char}$ is the chord $c$ and $u_{char}$ is the velocity magnitude at the exit boundary. For the time-step that is used in the computation, the cut-off Strouhal number $St_{\text{cutoff}}\approx 12600$. Figure 6 shows that for this probed point both Roe and HLLC produce similar trends of PSD for a wide range of the spectrum, but there are noticeable differences at very high frequencies. The reason for such differences can be related to the inherent numerical dissipation in both Riemann solvers and this would require further investigations.
@@ 4.3 High Pressure Turbine Vane with Heat Transfer
The third test case is the spanwise periodic ILES simulation of the linear cascade of a high pressure turbine vane LS89. The Mach number, Reynolds number, and temperature ratio between the inlet and exit are representative of typical industrial high-pressure turbines. The experimental data are available from Arts et al. [50]. The experiment involves studies of an uncooled turbine vane under a range of the Mach number, Reynolds number, and freestream turbulence. These conditions provide challenging and well-suited test cases for predicting the transition and heat transfer on the blade surface. There have been several successful LES/DNS simulations of this case using the high order finite-difference [51] method, finite-volume approach[52], and CPR [8]. The flow condition of the current study is MUR129. This case has a very low level of incoming turbulence ($0.8\%$) at the inlet. Previous work [8] prescribes no freestream turbulence at the inlet and obtains excellent agreement with the experimental data. The geometry detail and the flow condition for MUR129 that are used in this study are summarized in Table 5. The heat transfer coefficient (HTC) and the isentropic Mach number will be used to validate the current FR solver against the experimental data and to demonstrate the impact of different numerical configurations.
@@
The grid is generated by Gmsh. A quadrilateral mesh is generated in the blade-to-blade section and then extruded in the spanwise direction by $16.6\%$ of the chord. This value was used in the previous work by Jia and Wang [8] and shows a good prediction of HTC on the blade surface. Furthermore, Morata et al. [53] shows that the maximum difference of HTC on the surface of the blade is within $5\%$ when the span extrusion increases from $10\%$ to $20\%$ of the chord, but when the extrusion is reduced to $5\%$ of the chord, obvious differences in HTC are observed. Therefore, extrusion of the $16.6\%$ chord is a cost-effective option. Figure 9 shows the unstructured quadrilateral mesh in the blade-to-blade section and a close-up view of the mesh around the trailing edge. Several levels of spanwise grid spacing will be studied. In the coarsest one, 10 layers of hexahedra are extruded in the spanwise direction, and the total number of hexahedra is 244150. For the $p3$ simulation, the total number of degree of freedom (DOF) per equation is 15625600. The axial distance between the leading edge and the inlet boundary is approximately $c$. The axial distance between the trailing edge and the exit boundary is approximately $2c$. These distances are slightly larger than previous studies [8, 52] and this is beneficial to reduce the impact of the inflow/outflow boundaries on the solution. At the inlet, the total pressure, the total temperature, and two flow angles are specified. At the exit, a static pressure is specified. The upper and lower surfaces are set to be periodic.
@@
The computation starts with a $p_{0}$ run that efficiently creates a suitable initial flow field. Then the computation is restarted and the spatial precision order is increased to $p1$. The same procedures are then used to increase the order of accuracy from $p1$ to $p2$, and eventually from $p2$ to $p3$. A fixed CFL number of 0.85 is used in the computation. For different spatial orders of accuracy, this value is scaled by $\frac{1}{2p+1}$, where $p$ is the order of polynomials. For a $p3$ calculation, the average time step for HLLC-LMF is approximately $0.56\times 10^{-5}t_{c}$, which $t_{c}$ is the characteristic time.

### v2

@@ Abstract
A high-order flux reconstruction solver has been developed and validated to perform implicit large-eddy simulations of industrially representative turbomachinery flows. The T106c low-pressure turbine and VKI LS89 high-pressure turbine cases are studied. The solver uses the Rusanov Riemann solver to compute the inviscid fluxes on the wall boundaries, and HLLC or Roe to evaluate inviscid fluxes for internal faces. The impact of Riemann solvers is demonstrated in terms of accuracy and non-linear stability for turbomachinery flows. It is found that HLLC is more robust than Roe, but both Riemann solvers produce very similar results if stable solutions can be obtained. For non-linear stabilization, a local modal filter, which combines a smooth indicator and a modal filter, is used to stabilize the solution. This approach requires a tuning parameter for the smoothness criterion. Detailed analysis has been provided to guide the selection of a suitable value for different spatial orders of accuracy. This local-modal filter is also compared with the recent positivity-preserving entropy filter in terms of accuracy and stability for the LS89 turbine case. The entropy filter could stabilize the computation but is more dissipative than the local modal filter. Regarding the spanwise spacing of the grid, the case of the LS89 turbine shows that a $z^{+}$ of approximately $45-60$ is suitable for obtaining a satisfactory prediction of the heat transfer coefficient of the mean flow. This would allow for a coarse grid spacing in the spanwise direction and a cost-effective ILES aerothermal simulation for turbomachinery flows.
@@
discontinuous spectral element method , high order flux reconstruction , large-eddy simulation , turbomachinery
@@
[…]
@@
Effect of Riemann solvers on accuracy and numerical stability is investigated for turbomachinery flows.
@@
The positivity preserving entropy filter is compared with the local modal filter and assessed in turbomachinery flows.
@@ 1 Introduction
Computational Fluid Dynamics (CFD) has progressed significantly in the last few decades and has transformed the design process in the aerospace industry [1]. With the advancement of computing power and numerical algorithms, scale-resolving simulations, such as Large Eddy Simulation (LES), start complementing not only the experimental campaigns but also low-order numerical tools at the design stage. Reynolds-averaged Navier-Stokes (RANS) simulations and unsteady RANS (URANS) are the current workhorse in the industry for aerothermal design. However, it is well known that the RANS/URANS approach can face difficulties in reliably predicting vortex-dominated and separated flows and its associated heat transfer phenomenon.
The use of discontinuous spectral element methods (DSEM) has grown in prevalence in the past few years for LES due to their ability to achieve high-order spatial accuracy in unstructured grids and a compact stencil that is suited for massively parallel computations. Popular DSEM methods are the nodal Discontinuous Galerkin (DG) approach [2, 3] and, more recently, the Flux Reconstruction (FR) [4, 5] or the Correction Procedure via Reconstruction (CPR) [6] method. Despite the differences in these two methods, the numerical procedures developed in the collocation-based nodal DG could be applied to FR/CPR (such as filtering and boundary treatment) with little modification and vice versa. Implicit LES (ILES) is frequently used in DSEM, as the built-in dissipation of DSEM is found to be adequate as a subgrid model. The superior performance of DSEM over traditional second-order finite-volume solvers on ILES has been demonstrated by Vermeire et al. [7] and Jia and Wang [8].
@@
Filtering is an attractive stabilization technique due to its ease of implementation and computational efficiency. However, care must be taken when filtering is applied to DSEM, as it can significantly undermine the accuracy of DSEM [30]. The amount of filtering should be applied as much as needed to stabilize the solution but also as little as possible to avoid excessive dissipation. A modal filter could be localized using a smooth indicator so that the smooth-flow region is not subject to filtering. However, this strategy requires a smoothness criterion to determine whether the element should be filtered or not. This method has not been systematically studied in the literature for turbomachinery flows, and the choice of the smooth criterion is also not clear. The first contribution of this paper is: we will demonstrate the performance of this local modal filtering approach on turbomachinery flows and perform a detailed analysis of the smoothness criterion to provide a guide to select a suitable value. The recent entropy filter of Dzanic and Witherden [31] can also be considered as a localized modal filtering approach. Physical constraints are used as a “smoothness” criterion to determine which element should be filtered. This entropy filter is implemented and its performance is assessed in turbomachinery flows and compared with the local modal filter.
Riemann solvers have been found to have an impact on accuracy and numerical stability on simple turbulent flows [13] using DSEM. These effects have not been systematically studied for turbomachinery flows. In this work, such a study will be performed on industrially representative turbomachinery cases. In addition, Alhawwary and Wang [32] performed a study of the spanwise grid spacing on the prediction of mean flow variables (i.e., isentropic Mach number) in the T106c turbine using CPR. However, there have been no similar studies on the prediction of heat transfer coefficients in turbine blades, which is more challenging to predict than the pressure distribution on the blade surface. In this paper, we demonstrate this effect using the VKI LS89 turbine case. These are the second contribution of this work.
The computations in this paper are performed with a new in-house FR solver, AeroThermal High-Order Simulation (ATHOS). This paper is organized as follows. The governing equations of compressible flow and the FR method are briefly described. Then the local modal filter is introduced. The performance of the local modal filter on Taylor Green vortex, T106c, and LS89 turbine cases is demonstrated, the effect of Riemann solvers on accuracy and numerical stability is analyzed, and a detailed analysis of the smoothness criterion of the modal filter is presented. This is then followed by the conclusions and future work.
@@ 2.2 High-Order Flux Reconstruction
The FR method is an efficient discontinuous spectral element method to solve the partial differential equation (PDE) in differential form. It is briefly described here, and for a detailed description one can refer to Huynh [4], Vicent et al. [33] and Castonguay [34]. The NS equation of the compressible flow (Equation 1) is a second-order PDE. Before applying FR, it is first cast into a set of first order PDEs and this is shown in Equation 2.2 [29]:
@@ 2.3 Stabilization with filtering
DSEM methods can face robustness issues, and stabilization techniques are usually required to enhance robustness to handle underresolved flows or strong discontinuity (e.g., shocks). In this work, a local modal filtering (LMF) approach is developed. The general concept is to apply the modal filter in regions where stabilization is required while leaving the smooth-flow region untouched. Modal filtering is a simple and classical approach in DSEM and it can be represented as a matrix operation for the nodal solutions as [29]:
@@
where $\eta_{i}=\frac{i}{N_{p}-1}$, $i$ is the mode index and $N_{p}$ is the maximum order of the polynomial. $\eta_{c}$ is the cutoff mode threshold, below which the modes will not be filtered. $\alpha$ and $s$ determine the strength of the filter. $s$ is an even number and its values usually range from 2 to 8. $s$ is related to the scale selectivity of the filter and in practice a lower value of $s$ normally leads to a stronger filter. The value of $\alpha$ is $-\log_{10}(\varepsilon_{m})$ and $\varepsilon_{m}$ is the machine zero [29]. For simulations with double precision $\alpha\approx 36$. It is obvious that the power of the exponent in Equation 18 is a negative number; therefore, the modal filter is always dissipative.
@@ 4.1 Taylor Green Vortex at Re=1600 and Re=5000
Figure 2 shows the predicted time evolution of $\varepsilon_{K}$ using LMF at Reynolds numbers of 1600 and 5000, respectively. The results are compared with DNS data. Time $t$ is the normalized time, which is computed as $t=t_{\text{physical}}/t_{c}$. $t_{c}$ is the characteristic time and it is defined as $L/V_{0}$. At Re=1600, the DNS data is from Gassner and Beck [17]. For Re=5000 it is computed by Dairay et al. [45] using the Incompact3d code. From the comparison, it can be seen that there is no noticeable difference between the results of LMF and the one without filtering. Both results show excellent agreement with the reference DNS data. This confirms that LMF does not introduce noticeable dissipation to the simulation in the absence of strong discontinuity in the flow.
@@
The energy spectra is also examined at both Reynolds numbers and it is computed at $t=9$ and $t=14$, respectively. This is shown in Fig. 3. The spectra of the DNS data at Re=1600 is computed using a pseudo-spectral method on a $512^{3}$ mesh by Carton de Wiart et al. [46] and the one for Re=5000 is again computed by Dairay et al. [45]. The maximum resolvable wavenumber is marked by a vertical dash line. It can be seen that there is excellent agreement with the DNS data and there is no noticeable difference when LMF is used or not. It is noted that the purpose of this study is to demonstrate that when the flow is smooth, LMF essentially has a negligible effect on the flow field. But for more complicated flows, LMF will be switched on and this will be demonstrated in the following two turbine cases.
@@ 4.2 T106c Low Pressure Turbine
This test case performs the spanwise periodic ILES of the transitional and separated flow on the T106c turbine cascade. This test case has been well studied in the literature and is a standard test case in the International Workshop on High-Order CFD Methods. The flow in the T106c turbine is subsonic and is calculated at Re = 80,000. The Reynolds number is based on the blade chord and the flow conditions at the exit boundary. The inlet turbulence is very low and the flow features laminar separation and transition in the reattachment zone. The details of the geometry and boundary conditions are summarized in Table 3.
@@
The grid is generated by Gmsh [47] and the script is provided by the 4th International Workshop on High-Order CFD Methods  [^1] https://how4.cenaero.be/content/as2-spanwise-periodic-dnsles-transitional-turbine-cascades. A quadrilateral mesh is generated in the blade-to-blade section and a hexahedral mesh is generated by extruding the 2D quadrilateral mesh by $10\%$ of the chord for 12 layers. The script for the “baseline” mesh from the workshop is used and the only modification to the original script is that the expansion ratio of the boundary layer mesh is reduced from 1.3 to 1.2. The total number of hexahedral cells is 122904. For a $p3$ simulation, the total number of DOFs per equation is 7865856. The axial distance between the leading edge and the inlet boundary is approximately $1.6c$, where $c$ is the blade chord. The axial distance between the trailing edge and the exit boundary is approximately $2.4c$. Figure 4 shows the quadrilateral mesh in the blade-to-blade section and a close-up view of the boundary layer mesh around the trailing edge.
@@
Table 4 shows the numerical setup for the T106c turbine. LMF is used to stabilize the computation, HLLC and Roe Riemann solvers are used and compared. The values of $\alpha$ and $s$ are set to 36 and 6, respectively, and their values are fixed throughout this paper. According to Hesthaven and Warburton [29] both values provide a reasonable balance of numerical accuracy and stability. $s_{0}$ is the flow smoothness criterion and determines how ”local” the filter will be. If a large negative number is used, say $-16$, the filter will be applied to all elements. A value of -4 is used for both HLLC and Roe and successfully stabilizes the computation. More detailed analysis of the smooth criterion can be found in Section 4.4.1.
The computation starts from a $p_{0}$ run to efficiently wash away the initial transient flow field and creates a suitable initial flow field. Then the computation is restarted and the polynomial order is increased to $p1$. The same procedures are then used to increase the order of accuracy from $p1$ to $p2$, and eventually from $p2$ to $p3$. A fixed CFL number of 0.85 is used in the calculation. For different spatial orders of accuracy, this value is scaled by $\frac{1}{2p+1}$, where $p$ is the order of polynomial. For a $p3$ computation, the average time step is approximately $0.47\times 10^{-4}t_{c}$ for HLLC-LMF, where $t_{c}$ is the characteristic time and is defined as:
@@
Figure 5 shows the isosurfaces of the Q criterion for the instantaneous flow calculated by HLLC-LMF. Transitional flows on the suction side and vortex shedding can be clearly seen. Figure 6 shows the power spectral density (PSD) of the pressure at a point downstream of the trailing edge. The point is on the middle span plane and its coordinate on the blade-to-blade section is $(0.8591-0.5137)$. The frequency is represented in terms of the Strouhal number as:
@@
in which $f$ is the frequency, $L_{char}$ is the characteristic length and $u_{char}$. In this T106c case, $L_{char}$ is the chord $c$ and $u_{char}$ is the velocity magnitude at the exit boundary. For the time-step that is used in the computation, the cut-off Strouhal number $St_{\text{cutoff}}\approx 6300$. Figure 6 shows that for this probed point both Roe and HLLC agree with each other in the inertial sub-range, but there are noticeable differences at high frequencies. The reason for such differences can be related to the inherent numerical dissipation in both Riemann solvers and this would require further investigations.
@@ 4.3 High Pressure Turbine Vane with Heat Transfer
The third test case is the spanwise periodic ILES simulation of the linear cascade of a high-pressure turbine vane LS89. The Mach number, Reynolds number, and temperature ratio between the inlet and exit are representative of typical industrial high-pressure turbines. The experimental data are available from Arts et al. [50]. The experiment involves studies of an uncooled turbine vane under a range of the Mach number, Reynolds number, and freestream turbulence. These conditions provide challenging and well-suited test cases for predicting the transition and heat transfer on the blade surface. There have been several successful LES/DNS simulations of this case using the high-order finite-difference [51] method, finite-volume approach[52], and CPR [8]. The flow condition of the current study is MUR129. This case has a very low level of incoming turbulence ($0.8\%$) at the inlet. Previous work [8] prescribes no freestream turbulence at the inlet and obtains excellent agreement with the experimental data. The geometry detail and the flow condition for MUR129 that are used in this study are summarized in Table 5. The heat transfer coefficient (HTC) and the isentropic Mach number will be used to validate the current FR solver against the experimental data and to demonstrate the impact of different numerical configurations.
@@
The grid is generated by Gmsh. A quadrilateral mesh is generated in the blade-to-blade section and then extruded in the spanwise direction by $16.6\%$ of the chord. This value was used in the previous work by Jia and Wang [8] and shows a good prediction of HTC on the blade surface. Furthermore, Morata et al. [53] shows that the maximum difference of HTC on the blade surface is within $5\%$ when the span extrusion increases from $10\%$ to $20\%$ of the chord, but when the extrusion is reduced to $5\%$ of the chord, obvious differences in HTC are observed. Therefore, extrusion of the $16.6\%$ chord is a cost-effective option. Figure 9 shows the unstructured quadrilateral mesh in the blade-to-blade section and a close-up view of the mesh around the trailing edge. Several levels of spanwise grid spacing will be studied. In the coarsest one, 10 layers of hexahedra are extruded in the spanwise direction, and the total number of hexahedra is 244150. For the $p3$ simulation, the total number of degrees of freedom (DOF) per equation is 15625600. The axial distance between the leading edge and the inlet boundary is approximately $c$. The axial distance between the trailing edge and the exit boundary is approximately $2c$. These distances are slightly larger than previous studies [8, 52] and this is beneficial to reduce the impact of the inflow/outflow boundaries on the solution. At the inlet, the total pressure, the total temperature and two flow angles are specified. At the exit, a static pressure is specified. The upper and lower surfaces are set to be periodic.
@@
The computation starts with a $p_{0}$ run that efficiently creates a suitable initial flow field. Then the computation is restarted and the spatial precision order is increased to $p1$. The same procedures are then used to increase the order of accuracy from $p1$ to $p2$, and eventually from $p2$ to $p3$. A fixed CFL number of 0.85 is used in the computation. For different spatial order of accuracy, this value is scaled by $\frac{1}{2p+1}$, where $p$ is the order of polynomials. For a $p3$ calculation, the average time step for HLLC-LMF is approximately $0.56\times 10^{-5}t_{c}$, which $t_{c}$ is the characteristic time.